### v3

@@ 6.1 Significance of the results
Figure 14: A diagram illustrating the structural evolution we’ve proposed for the Milky Way. The gray ellipse shows the DM halo, with the darker region representing the inner halo associated with the stellar disk. The black dots denote satellite galaxies, while the red curve represents orbit of a massive satellite retrograde to the rotation of stellar disk. Several Gyr ago, the Milky Way featured an oblate DM halo, with the stellar disk, DM halo, and satellite galaxies co-aligned. Around 6-8 Gyr ago, the disk and inner halo began tilting, and this tilt continued over the subsequent Gyr, eventually leading to a twisted DM halo, with the outer DM halo and satellite plane becoming perpendicular to the stellar disk.
@@
In our default model with a constant axis ratio $q_{\rm DM}=Z/X$ and $p_{\rm DM}=Y/X$, the axis ratio $q_{\rm DM}$ is constrained to be $0.92\pm 0.08$, consistent with most previous studies (e.g., Wegg2019MNRAS.485.3296W; Huang2024NatAs...8.1294H; Zhang2025). However, we have also shown that the Milky Way DM halo is triaxial, with $q_{\rm DM}>p_{\rm DM}$ strongly preferred. Although there is still some degeneracy between the two parameters, with relatively large uncertainty on the axis ratio $p_{\rm DM}$ from our dynamical model, the allowed region in $q_{\rm DM}$ vs. $p_{\rm DM}$ is already narrow enough for us to distinguish the galaxy formation scenarios. Except for the four Milky Way analogues aforementioned, TNG50 501208, TNG50 546474, TNG50 483594, and Auriga 12, all the other TNG50 galaxies fall within the $3\sigma$ confidence level of our model, displaying nearly isotropic or coplanar satellite distributions, as expected from their intrinsic near-spherical or normally oblate DM halos (see Appendix B).
@@
The presence of a vertical satellite plane thus helps to pin down the shape of the DM halo: a nearly oblate but vertically orientated DM halo is the only type allowed within the $3\sigma$ threshold of the Milky Way that is associated with a vertical satellite plane. In three of the four galaxies, the vertical alignment of the DM halo and satellite plane is purely due to the identical physical mechanism: a flip of the stellar disk, likely torqued by minor mergers that spiralled inward. They are also the only three galaxies with DM halo shapes lying within $1\sigma$ range of the Milky Way. In the remaining case, TNG50 483594, its vertical alignment was initialised by variations in the filament network along with mass accreted into the galaxy, but was strengthened by disk tilt. Its DM halo is intrinsically rounder, out of $1\sigma$ range of the Milky Way. What’s more, it has a very different assembly history from the Milky Way. We therefore strongly propose that the Milky Way arises from the same scenario as the three $1\sigma$ galaxies; otherwise, a system combining such a DM halo shape and vertical satellite plane would be highly impossible in either the TNG50 or Auriga simulations.
@@
The merger history of the Milky Way closely resembles that of two of the aforementioned $1\sigma$ galaxies, TNG50 501208 and TNG50 546474. It is believed that the stellar disk formed after the last major merger (Belokurov2020MNRAS.494.3880B)–the Gaia-Enceladus-Sausage (GSE) merger occurred 8-10 Gyr ago (Helmi2018Natur.563...85H; Belokurov2018MNRAS.478..611B), comparable to the major mergers that happened in these two galaxies 9-10 Gyr ago (highlighted in Fig 12). The prominent Sagittarius stream (Majewski2003ApJ...599.1082M; Belokurov2006ApJ...642L.137B) imprints several spiral-in orbital passages in the halo over the past 6-8 Gyr (Dierickx2017ApJ...836...92D). Recent studies suggest that the progenitor mass of Sagittarius is likely to be $5-6\times 10^{10}$M  or even larger (Gibbons2017MNRAS.464..794G; Read2019MNRAS.487.5799R), comparable to the minor mergers that induced the disk flip in TNG50 501208 and TNG50 546474.
@@
Direct N-body simulations (Dodge2023MNRAS.518.2870D) show that the GSE could effectively tilt the Galactic disk, if not destroy it. Although GSE-like major mergers usually result in the co-alignment of the disk and DM halo (Dillamore2022MNRAS.513.1867D), the ultimate orientation of the disk depends on the orbital properties of the satellite. We find that satellites coming in on prograde orbits usually result in the co-alignment of the disk and orbital plane, while satellites coming in on retrograde orbits could result in disk vertical to the satellite orbital plane (Cai & Zhu, in preparation). The latter is consistent with the case in the MW that GSE remnants are with retrograde rotations (Helmi2018Natur.563...85H). We suggest that Sagittarius could also be a promising candidate that induced the disk tilting to vertical alignment, similar to the tilt that began about 6 Gyr ago in TNG50 501208. In real cases, the torques of multiple mergers could have worked together, along with the accretion of unbound DM (Genel2010ApJ...719..229G).
@@
A diagram is included in Fig.14 to illustrate the structural evolution we propose for the Milky Way. The current disk-halo orientation of the Milky Way may not be fully stable; the disk could still be tilting. Although there are currently no direct observational constraints, the model of the stream Pal 5’s tidal tails favours a tilting rate of $15^{\circ}$ per Gyr for the Milky Way disk (Nibauer2024ApJ...969...55N), supporting our scenario.
6.2 The effects of a tilting disk on dynamical models
@@
If the Galactic disk is still tilting at $z=0$, a tilting disk might affect our dynamical models in two aspects: first, the overall potential is evolving, and thus the stellar halo, especially the inner halo, is not necessarily in a stationary state; second, the inner halo may be tilting with the disk and cause a velocity shift and finally, a twist compared to the outer halo, as suggested by the simulations. We will evaluate the significance of these effects, considering a disk tilt rate of $15^{\circ}$ per Gyr.
@@
The cross time of stars in the inner stellar halo is 0.1-0.2 Gyr. During one period of cross time, the disk could tilt about two degrees, which has negligible effects on the potential. In other words, the potential evolves such that the motion of stars will be affected, but the evolution is so slow that the halo stars can update their motions efficiently to reach dynamical equilibrium in a timely manner. Our assumption of dynamical equilibrium should still hold with any momentum. Moreover, the dynamical equilibrium of the system is indeed supported by the good match between the data and our best-fitting model.
@@
If there is a velocity shift between the inner and outer halo caused by the tilt of the inner halo with the disk, it will lead to some non-equilibrium features in the outer halo. A disk tilt rate of $15^{\circ}$ per Gyr equals 0.05 mas/yr, and corresponds to 8 km/s at $r=30$ kpc and 13 km/s at $r=50$ kpc, which are less than the net velocities we subtracted for the correction of LMC effects. The effects of tilt to the outer halo, if they exist, will be mixed with the LMC effects we subtracted. As we tested, imposing the LMC correction or not does not make a noticeable difference in our final results. We thus do not expect a noticeable difference in our final results if there is a tilt of the inner halo with the disk.
In summary, the disk tilt, if it is still ongoing at $z=0$, is actually very slow and has minor effects on the kinematics of the stellar halo, which should not affect the results of the dynamical models.
@@ 6.3 The tilt angle of the dark matter halo
We fixed the tilt angle of the DM halo $\beta_{q}$ to be either $0^{\circ}$ or $90^{\circ}$ compared to the disk, by fixing the Z axis as perpendicular to the stellar disk. Our current approach allows for a fair comparison with the cosmological simulations by defining exactly the same coordinates for measuring the DM shape of the simulated galaxies. In reality, the tilt angle of the DM halo is not likely to be exactly $0^{\circ}$ or $90^{\circ}$; a disk vertical to the intermediate axis of the DM halo was suggested to be unstable, while a disk can persist off one of the principal planes of the potential (Debattista2013MNRAS.434.2971D). However, the disk tilt angle is strongly degenerated with the axis ratio $q_{\rm DM}$ for the measurement. A tilt angle of $\beta_{q}\sim 15^{\circ}$ of the DM halo compared to the stellar disk was preferred in creating the disk warp (Han2023NatAs...7.1481H) and the presence of a tilted stellar halo dominated by GSE (Han2022ApJ...934...14H). The tilted stellar halo was used to constrain an orbit-superposition model (Dillamore2025arXiv251000095D), in which the DM tilt angle is allowed to be free. They obtain an average tilt angle of $\beta_{q}\sim 45^{\circ}$ for the DM halo, but there is a strong degeneracy between the two parameters such that $\beta_{q}\sim 30-80^{\circ}$ and $q_{\rm DM}\sim 0.7-0.95$ are allowed within their $1\sigma$ confidence level. Note that all the above results are mainly constrained by GSE at $r\lesssim 30$ kpc. A tilted DM halo in the inner regions is generally consistent with our results; we find tentative evidence for a radially-varying shape/orientation of the DM halo, which being less tilted in the inner regions and becomes vertically aligned in the outer regions, consistent with Vasiliev2021MNRAS.501.2279V. Data with larger spatial coverage are needed to further constrain the twist of the DM halo. Interestingly, A recent study has constructed the 3D distribution of stellar halo out to about 50 kpc using Desi observations, which found the Milky Way stellar halo is twisted and becomes vertical to the stellar disk in the outer regions (Li et al., in prep), which turns to be consistent with the shape of the DM halo we found.
@@ 7 Summary
We determine the 3D DM distribution of the Milky Way by utilizing the empirical triaxial orbit-superposition model applied to the 6D phase-space data of halo stars provided by LAMOST and Gaia. This model accommodates a highly flexible triaxial DM halo, where the axis ratios may change with varying radii. Both the radial density profile and the 3D shape of the DM halo are constrained simultaneously. Furthermore, we benchmark our findings against 105 galaxies resembling the Milky Way from cosmological simulations. Our major findings are as follows.
  • We discover a triaxial, nearly oblate DM halo with $q_{\rm DM}=Z/X=0.92\pm 0.08$, $p=Y/X=0.8\pm 0.2$ averaging $r\lesssim 50$ kpc, where the $Z$ axis is defined perpendicular to the stellar disk. The axes ratio $q_{\rm DM}>p_{\rm DM}$ is strongly preferred; the long-intermediate axis plane of the DM halo is unexpectedly vertical to the Galactic disk.
@@
  • This striking configuration of the DM halo and satellite system strongly suggests that the Galactic disk has flipped, torqued by minor mergers, from its original alignment with the DM halo and satellite plane, as evidenced by Milky Way analogues we found in Auriga and TNG50.
  • We find tentative evidence that the Milky Way DM halo is twisted, consistent with alignment with the disk in the inner $r\lesssim 20$ kpc, and becomes vertically orientated in the outer regions, consistent with the prediction of disk flip scenario.
@@ Acknowledgements.
The authors thank David Hogg, Jianhui Lian, Zhaozhou Li, Wenting Wang for useful discussions. The authors acknowledge the support from the CAS Project for Young Scientists in Basic Research under grant No. YSBR-062 (LZ, X.X.X, LZhang), the National Key R&D Program of China No. 2025YFF0511002(LZ), 2022YFF0503403 (LZ), 2022YFA1602903 (XK), and No. 2024YFA1611902 (X.X.X), and the science research grants from the China Manned Space Project with NO. CMS-CSST-2025-A11 (X.X.X, LZ, CY, LZhang), the Strategic Priority Research Program of Chinese Academy of Sciences grant No. XDB1160102 (X.X.X), and National Natural Science Foundation of China (NSFC) No. 12588202 (X.X.X).
@@ Appendix A Quantifying the spatial arrangement of satellites
First, we compute their angular momentum; the orientation of the angular momentum $\theta_{L}$ is determined by the angle between the spin direction and its projection line on the disk plane. A satellite with $\theta_{L}\approx 0$ is in an orbit vertically aligned with the stellar disk. The $f_{\rm Sat,vertical}$ is defined as the fraction of satellites with $\theta_{L}<45^{\circ}$. We evaluated the azimuthal anisotropy of the satellite distribution by calculating their azimuthal angle $\phi_{r}$ in the disk plane. The distribution of all satellites’ $\phi_{r}$ is plotted, and we fit the histogram with a sine function $n=n_{1}\sin(2\phi_{r}+\phi_{0})+n_{0}$. An anisotropy parameter, $\xi_{\rm Sat,\phi}=n_{1}/n_{0}$, is defined. The definition of these two parameters is illustrated in Fig. 15. Under this definition, a perfectly isotropic satellite system has $f_{\rm Sat,vertical}=0.707$ and $\xi_{\rm Sat,\phi}=0$.
For the Milky Way, we took 52 satellite galaxies with measured 3D position and velocity (Pace2022ApJ...940..136P), and added LMC to the data, thus with 53 in total. As shown in Figure 15, the satellite galaxies of the Milky Way are vertically orbiting with $f_{\rm Sat,vertical}=0.87$ and tend to align in a plane with high azimuthal anisotropy $\xi_{\rm Sat,\phi}=0.55$.
@@
The results of these 94 TNG50 and 11 Auriga galaxies are presented in Fig.16. As shown in the figure, the distribution of satellite galaxies around most galaxies is not entirely isotropic. Note that these parameters are affected by the number of satellites and there are still certain scatters in describing the satellite arrangement. However, we notice that there is a general correlation between the shape of the DM halo and the arrangement of the satellite system; the correlation is stronger for the DM shape measured at a larger radius. In galaxies with oblate DM halos ($q_{\rm DM}<p_{\rm DM}$), satellite galaxies tend to align coplanarly with the stellar disk, characterised by $f_{\rm Sat,vertical}<0.707$ and $\xi_{\rm Sat,\phi}\lesssim 0.3$. This configuration is typical for systems with non-zero total angular momentum, where the orientation of the stellar disk, DM halo, and satellite system are horizontally synchronised, consistent with previous results (Shao2016MNRAS.460.3772S). There are a number of galaxies with satellites distributed highly coplanar with $f_{\rm Sat,vertical}<0.5$; a plane of satellites like the Milky Way (not considering its orientation) is thus not so rare, as also revealed by previous studies (Sawala2023NatAs...7..481S). However, galaxies with triaxial or near-spherical DM halos ($q_{\rm DM}\simeq p_{\rm DM}$) show a more isotropic distribution of satellites ($f_{\rm Sat,vertical}\sim 0.707$). Meanwhile, in the rare cases of galaxies with DM halo vertically aligned with the stellar disk ($q_{\rm DM}>p_{\rm DM}$), their satellite systems appear vertically orientated $f_{\rm Sat,vertical}>0.707$, and at the same time show considerable anisotropy in the azimuthal direction (large value of $\xi_{\rm Sat,\phi}$).
@@
Figure 16: Arrangement of satellite system in the Milky Way and the Milky Way-like simulated galaxies, and relation with the 3D shape of DM halo. Left: the configuration of satellite system of each galaxy is characterized by two parameters: the fraction of satellites moving on vertical orbits $f_{\rm Sats,vertical}$, and the azimuthal anisotropy of the satellite distribution $\xi_{\rm Sat,\phi}$. Each dot is one galaxy colored by their DM halo shape $q_{\rm DM}-p_{\rm DM}$ (<50<50 kpc). The red star represents the Milky Way. The 13 TNG50/Auriga galaxies having DM halo shapes within $3\sigma$ confidence level of the Milky Way are highlighted. Right: correlation of the DM halo shape $q_{\rm DM}-p_{\rm DM}$ versus satellite distribution $f_{\rm Sats,vertical}$ of the 94 TNG50 galaxies, $\mathcal{R}$ in the legend is the Pearson correlation coefficiant. The correlation becomes stronger when increasing the number of satellites by taking a lower mass threshold and measuring DM shape at a larger radius.
Figure 17: Orbital configuration of satellites in the Milky Way and TNG50/Auriga galaxies within the $3\sigma$ confidence level. For each galaxy, we show the orbits of the ten brightest/most massive satellites. For the Milky Way, the pluses represent the orbit of the Sagittarius stream. Colors represent the relative mass of the satellites in each galaxy, red is the most massive one; the thickness of the curves are scaled with the maximum history mass of the satellites. The blue ellipses in the top panels indicate the orientation of stellar disk. The first four simulated galaxies in the top-left panels are the four galaxies we discussed, they show similar vertical satellite plane as the the Milky Way. While the rest nine simulated galaxies show rather isotropic distribution of the satellite galaxies. Eight of the 13 galaxies (highlighted by red boxes), experienced disk tilt from $30^{\circ}$ to $100^{\circ}$, most of them are likely induced by a massive minor merger (red curves), mostly spiraled inward on tangential orbits, except 580406.

### v4

@@ 6.1 Significance of the results
Figure 14: Diagram illustrating the structural evolution that we have proposed for the Milky Way. The grey ellipse shows the DM halo, with the darker region representing the inner halo associated with the stellar disc. The black dots denote satellite galaxies, while the red curve represents orbit of a massive satellite retrograde to the rotation of stellar disc. Several gigayears ago, the Milky Way featured an oblate DM halo, with the stellar disc, DM halo, and satellite galaxies co-aligned. Around 6-8 Gyr ago, the disc and inner halo began tilting, and this tilt continued over the subsequent gigayears, eventually leading to a twisted DM halo, with the outer DM halo and satellite plane becoming perpendicular to the stellar disc.
@@
In our default model with a constant axis ratio of $q_{\rm DM}=Z/X$ and $p_{\rm DM}=Y/X$, the axis ratio $q_{\rm DM}$ is constrained to be $0.92\pm 0.08$, consistent with most previous studies (e.g., Wegg et al., 2019; Huang et al., 2024; Zhang et al., 2025). However, we have also shown that the Milky Way DM halo is triaxial, with $q_{\rm DM}>p_{\rm DM}$ strongly preferred. Although there is still some degeneracy between the two parameters, with a relatively large uncertainty on the axis ratio $p_{\rm DM}$ from our dynamical model, the allowed region in $q_{\rm DM}$ versus $p_{\rm DM}$ is already narrow enough for us to distinguish the galaxy formation scenarios. Except for the four Milky Way analogues aforementioned, TNG50 501208, TNG50 546474, TNG50 483594, and Auriga 12, all the other TNG50 galaxies fall within the $3\sigma$ confidence level of our model, displaying nearly isotropic or coplanar satellite distributions, as was expected from their intrinsic near-spherical or normally oblate DM halos (see Appendix B).
@@
The presence of a vertical satellite plane thus helps to pin down the shape of the DM halo: a nearly oblate but vertically orientated DM halo is the only type allowed within the $3\sigma$ threshold of the Milky Way that is associated with a vertical satellite plane. In three of the four galaxies, the vertical alignment of the DM halo and satellite plane is purely due to the identical physical mechanism: a flip of the stellar disc, likely torqued by minor mergers that spiralled inwards. They are also the only three galaxies with DM halo shapes lying within $1\sigma$ range of the Milky Way. In the remaining case, TNG50 483594, its vertical alignment was initialised by variations in the filament network along with mass accreted into the galaxy, but was strengthened by disc tilt. Its DM halo is intrinsically rounder, out of $1\sigma$ range of the Milky Way. What is more, it has a very different assembly history from the Milky Way. We therefore strongly propose that the Milky Way arises from the same scenario as the three $1\sigma$ galaxies; otherwise, a system combining such a DM halo shape and vertical satellite plane would be highly impossible in either the TNG50 or Auriga simulations.
@@
The merger history of the Milky Way closely resembles that of two of the aforementioned $1\sigma$ galaxies, TNG50 501208 and TNG50 546474. It is believed that the stellar disc formed after the last major merger (Belokurov et al., 2020) – the Gaia-Enceladus-Sausage (GSE) merger – occurred 8-10 Gyr ago (Helmi et al., 2018; Belokurov et al., 2018), comparable to the major mergers that happened in these two galaxies 9-10 Gyr ago (highlighted in Fig 12). The prominent Sagittarius stream (Majewski et al., 2003; Belokurov et al., 2006) imprints several spiral-in orbital passages in the halo over the past 6-8 Gyr (Dierickx and Loeb, 2017). Recent studies suggest that the progenitor mass of Sagittarius is likely to be $5-6\times 10^{10}$M  or even larger (Gibbons et al., 2017; Read and Erkal, 2019), comparable to the minor mergers that induced the disc flip in TNG50 501208 and TNG50 546474.
@@
Direct N-body simulations (Dodge et al., 2023) show that the GSE could effectively tilt the Galactic disc, if not destroy it. Although GSE-like major mergers usually result in the co-alignment of the disc and DM halo (Dillamore et al., 2022), the ultimate orientation of the disc depends on the orbital properties of the satellite. We find that satellites coming in on prograde orbits usually result in the co-alignment of the disc and orbital plane, while satellites coming in on retrograde orbits could result in a disc vertical to the satellite orbital plane (Cai & Zhu et al., in preparation). The latter is consistent with the case in the MW that GSE remnants are with retrograde rotations (Helmi et al., 2018). We suggest that Sagittarius could also be a promising candidate that induced the disc tilting to vertical alignment, similar to the tilt that began about 6 Gyr ago in TNG50 501208. In real cases, the torques of multiple mergers could have worked together, along with the accretion of unbound DM (Genel et al., 2010).
@@
A diagram is included in Fig.14 to illustrate the structural evolution we propose for the Milky Way. The current disc-halo orientation of the Milky Way may not be fully stable; the disc could still be tilting. Although there are currently no direct observational constraints, the model of the stream Pal 5’s tidal tails favours a tilting rate of $15^{\circ}$ per Gyr for the Milky Way disc (Nibauer et al., 2024), supporting our scenario.
6.2 The effects of a tilting disc on dynamical models
@@
If the Galactic disc is still tilting at $z=0$, a tilting disc might affect our dynamical models in two aspects: first, the overall potential is evolving, and thus the stellar halo, especially the inner halo, is not necessarily in a stationary state; second, the inner halo may be tilting with the disc and cause a velocity shift and finally, a twist compared to the outer halo, as suggested by the simulations. We evaluate the significance of these effects, considering a disc tilt rate of $15^{\circ}$ per Gyr.
@@
The cross time of stars in the inner stellar halo is 0.1-0.2 Gyr. During one period of cross time, the disc could tilt about two degrees, which has negligible effects on the potential. In other words, the potential evolves such that the motion of stars will be affected, but the evolution is so slow that the halo stars can update their motions efficiently to reach dynamical equilibrium in a timely manner. Our assumption of dynamical equilibrium should still hold with any momentum. Moreover, the dynamical equilibrium of the system is indeed supported by the good match between the data and our best-fitting model.
@@
If there is a velocity shift between the inner and outer halo caused by the tilt of the inner halo with the disc, it will lead to some non-equilibrium features in the outer halo. A disc tilt rate of $15^{\circ}$ per Gyr equals 0.05 mas/yr, and corresponds to 8 km/s at $r=30$ kpc and 13 km/s at $r=50$ kpc, which are less than the net velocities we subtracted for the correction of LMC effects. The effects of tilt to the outer halo, if they exist, will be mixed with the LMC effects we subtracted. As we tested, imposing the LMC correction or not does not make a noticeable difference in our final results. We thus do not expect a noticeable difference in our final results if there is a tilt of the inner halo with the disc. In summary, the disc tilt, if it is still ongoing at $z=0$, is actually very slow and has minor effects on the kinematics of the stellar halo, which should not affect the results of the dynamical models.
@@ 6.3 The tilt angle of the dark matter halo
We fixed the tilt angle of the DM halo, $\beta_{q}$, to be either $0^{\circ}$ or $90^{\circ}$ compared to the disc, by fixing the Z axis as perpendicular to the stellar disc. Our current approach allows for a fair comparison with the cosmological simulations by defining exactly the same co-ordinates for measuring the DM shape of the simulated galaxies. In reality, the tilt angle of the DM halo is not likely to be exactly $0^{\circ}$ or $90^{\circ}$; a disc vertical to the intermediate axis of the DM halo was suggested to be unstable, while a disc can persist off one of the principal planes of the potential (Debattista et al., 2013). However, the disc tilt angle is strongly degenerated with the axis ratio $q_{\rm DM}$ for the measurement. A tilt angle of $\beta_{q}\sim 15^{\circ}$ of the DM halo compared to the stellar disc was preferred in creating the disc warp (Han et al., 2023) and the presence of a tilted stellar halo dominated by GSE (Han et al., 2022b). The tilted stellar halo was used to constrain an orbit-superposition model (Dillamore and Sanders, 2025), in which the DM tilt angle is allowed to be free. They obtain an average tilt angle of $\beta_{q}\sim 45^{\circ}$ for the DM halo, but there is a strong degeneracy between the two parameters such that $\beta_{q}\sim 30-80^{\circ}$ and $q_{\rm DM}\sim 0.7-0.95$ are allowed within their $1\sigma$ confidence level. Note that all the above results are mainly constrained by GSE at $r\lesssim 30$ kpc. A tilted DM halo in the inner regions is generally consistent with our results; we find tentative evidence for a radially varying shape of the DM halo, which being less tilted in the inner regions and becomes vertically aligned in the outer regions, consistent with Vasiliev et al. (2021). Data with larger spatial coverage are needed to further constrain the twist of the DM halo. Interestingly, a recent study has constructed the 3D distribution of the stellar halo out to about 50 kpc using Desi observations, which found that the Milky Way stellar halo is twisted and becomes vertical to the stellar disc in the outer regions (Li2025arXiv251201350L), which turns out to be consistent with the shape of the DM halo we found.
@@ 7 Summary
We determined the 3D DM distribution of the Milky Way by utilising the empirical triaxial orbit-superposition model applied to the 6D phase-space data of halo stars provided by LAMOST and Gaia. This model accommodates a highly flexible triaxial DM halo, where the axis ratios may change with varying radii. Both the radial density profile and the 3D shape of the DM halo are constrained simultaneously. Furthermore, we benchmark our findings against 105 galaxies resembling the Milky Way from cosmological simulations. Our major findings are as follows.
  • We discover a triaxial, nearly oblate DM halo with $q_{\rm DM}=Z/X=0.92\pm 0.08$, $p=Y/X=0.8\pm 0.2$ averaging $r\lesssim 50$ kpc, where the $Z$ axis is defined perpendicular to the stellar disc. The axis ratio $q_{\rm DM}>p_{\rm DM}$ is strongly preferred; the long-intermediate axis plane of the DM halo is unexpectedly vertical to the Galactic disc.
@@
  • This striking configuration of the DM halo and satellite system strongly suggests that the Galactic disc has flipped, torqued by minor mergers, from its original alignment with the DM halo and satellite plane, as is evidenced by Milky Way analogues we found in Auriga and TNG50.
  • We find tentative evidence that the Milky Way DM halo is twisted, consistent with alignment with the disc in the inner $r\lesssim 20$ kpc, and becomes vertically orientated in the outer regions, consistent with the prediction of disc flip scenario.
@@ Acknowledgements.
The authors thank David Hogg, Jianhui Lian, Zhaozhou Li, Wenting Wang for useful discussions. The authors acknowledge the support from the National Key R&D Program of China No. 2025YFF0511002(LZ), 2022YFF0503403 (LZ), 2022YFA1602903 (XK), and No. 2024YFA1611902 (X.X.X), the CAS Project for Young Scientists in Basic Research under grant No. YSBR-062 (LZ, X.X.X, LZhang), and the science research grants from the China Manned Space Project with NO. CMS-CSST-2025-A11 (X.X.X, LZ, CY, LZhang), the Strategic Priority Research Program of Chinese Academy of Sciences grant No. XDB1160102 (X.X.X), and National Natural Science Foundation of China (NSFC) No. 12588202 (X.X.X).
@@ Appendix A Quantifying the spatial arrangement of satellites
First, we compute their angular momentum; the orientation of the angular momentum $\theta_{L}$ is determined by the angle between the spin direction and its projection line on the disc plane. A satellite with $\theta_{L}\approx 0$ is in an orbit vertically aligned with the stellar disc. The $f_{\rm Sat,vertical}$ is defined as the fraction of satellites with $\theta_{L}<45^{\circ}$. We evaluated the azimuthal anisotropy of the satellite distribution by calculating their azimuthal angle $\phi_{r}$ in the disc plane. The distribution of all satellites’ $\phi_{r}$ is plotted, and we fit the histogram with a sine function $n=n_{1}\sin(2\phi_{r}+\phi_{0})+n_{0}$. An anisotropy parameter, $\xi_{\rm Sat,\phi}=n_{1}/n_{0}$, is defined. The definition of these two parameters is illustrated in Fig. 15. Under this definition, a perfectly isotropic satellite system has $f_{\rm Sat,vertical}=0.707$ and $\xi_{\rm Sat,\phi}=0$.
For the Milky Way, we took 52 satellite galaxies with measured 3D position and velocity (Pace et al. 2022), and added LMC to the data, thus with 53 in total. As shown in Figure 15, the satellite galaxies of the Milky Way are vertically orbiting with $f_{\rm Sat,vertical}=0.87$ and tend to align in a plane with high azimuthal anisotropy $\xi_{\rm Sat,\phi}=0.55$.
@@
The results of these 94 TNG50 and 11 Auriga galaxies are presented in Fig.16. As shown in the figure, the distribution of satellite galaxies around most galaxies is not entirely isotropic. Note that these parameters are affected by the number of satellites and there are still certain scatters in describing the satellite arrangement. However, we notice that there is a general correlation between the shape of the DM halo and the arrangement of the satellite system; the correlation is stronger for the DM shape measured at a larger radius. In galaxies with oblate DM halos ($q_{\rm DM}<p_{\rm DM}$), satellite galaxies tend to align coplanarly with the stellar disc, characterised by $f_{\rm Sat,vertical}<0.707$ and $\xi_{\rm Sat,\phi}\lesssim 0.3$. This configuration is typical for systems with non-zero total angular momentum, where the orientation of the stellar disc, DM halo, and satellite system are horizontally synchronised, consistent with previous results (Shao et al. 2016). There are a number of galaxies with satellites distributed highly coplanar with $f_{\rm Sat,vertical}<0.5$; a plane of satellites like the Milky Way (not considering its orientation) is thus not so rare, as also revealed by previous studies (Sawala et al. 2023). However, galaxies with triaxial or near-spherical DM halos ($q_{\rm DM}\simeq p_{\rm DM}$) show a more isotropic distribution of satellites ($f_{\rm Sat,vertical}\sim 0.707$). Meanwhile, in the rare cases of galaxies with DM halo vertically aligned with the stellar disc ($q_{\rm DM}>p_{\rm DM}$), their satellite systems appear vertically orientated $f_{\rm Sat,vertical}>0.707$, and at the same time show considerable anisotropy in the azimuthal direction (large value of $\xi_{\rm Sat,\phi}$).
@@
Figure 16: Arrangement of satellite system in the Milky Way and the Milky Way-like simulated galaxies, and relation with the 3D shape of DM halo. Left: the configuration of satellite system of each galaxy is characterised by two parameters: the fraction of satellites moving on vertical orbits $f_{\rm Sats,vertical}$, and the azimuthal anisotropy of the satellite distribution $\xi_{\rm Sat,\phi}$. Each dot is one galaxy coloured by their DM halo shape $q_{\rm DM}-p_{\rm DM}$ (<50<50 kpc). The red star represents the Milky Way. The 13 TNG50/Auriga galaxies having DM halo shapes within $3\sigma$ confidence level of the Milky Way are highlighted. Right: correlation of the DM halo shape $q_{\rm DM}-p_{\rm DM}$ versus satellite distribution $f_{\rm Sats,vertical}$ of the 94 TNG50 galaxies, $\mathcal{R}$ in the legend is the Pearson correlation coefficiant. The correlation becomes stronger when increasing the number of satellites by taking a lower mass threshold and measuring DM shape at a larger radius.
Figure 17: Orbital configuration of satellites in the Milky Way and TNG50/Auriga galaxies within the $3\sigma$ confidence level. For each galaxy, we show the orbits of the ten brightest/most massive satellites. For the Milky Way, the pluses represent the orbit of the Sagittarius stream. The colours represent the relative mass of the satellites in each galaxy, red is the most massive one; the thickness of the curves are scaled with the maximum history mass of the satellites. The blue ellipses in the top panels indicate the orientation of stellar disc. The first four simulated galaxies in the top-left panels are the four galaxies we discussed, they show similar vertical satellite plane as the the Milky Way. While the rest nine simulated galaxies show rather isotropic distribution of the satellite galaxies. Eight of the 13 galaxies (highlighted by red boxes), experienced disc tilt from $30^{\circ}$ to $100^{\circ}$, most of them are likely induced by a massive minor merger (red curves), mostly spiraled inward on tangential orbits, except 580406.